Chris is 3 years older than Jacob. Nathan is eighteen years old. Two years ago the sum of all their ages was 59 years old. How old is Jacob?

Answer by rm29924(97)

Let x = Jacob's current age
Let x+3 = Chris's current age
18 = Nathan's current age
Jacob's age 2 years ago is x-2
Chris's age 2 years ago is x+3-2 which is x+1
Nathan's age 2 years ago is 18-2 which is 16
two years ago all their ages added up to 59 years old
so our equation is:
x-2+x+1+16=59
2x+15=59
2x=44
x=22
Jacob is 22 years old
plug the answer into the equation to check work